OpenStudy (anonymous):

Use the quadratic formula to find the zeros of the function. y = 16x2 + 40x + 25 A. {-1.25, 1.25} B. {-1.25} C. {-4, 5} D. {-0.8} Use the quadratic formula to find the zeros of the function. y = 16x2 + 40x + 25 A. {-1.25, 1.25} B. {-1.25} C. {-4, 5} D. {-0.8} @Mathematics

OpenStudy (anonymous):

To find the roots ("zeros") of a quadratic, the quadratic formula:\[\left( -b \pm\sqrt{{b^2}-(4ac))}\right) \div 2a\]\[=\left( -40 \pm \sqrt{{40^2}-(4 \times 16 \times 25)} \right) \div (2 \times 16)\]\[=\left(-40 \pm \sqrt{1600 - 1600} \right) \div 32\]\[=-40 \div 32\]\[=-1.25\] Often there will be two roots, given that the plus/minus will generally lead to two different answers.